# A Polylogarithmic Competitive Algorithm for Stochastic Online Sorting and TSP

### v2

@@ 16    $x\leftarrow\texttt{receive\_sample}(\mathcal{D})$;
17    $k_{1}\leftarrow\texttt{index}(\mathcal{T},i-1)$, $k_{2}\leftarrow\texttt{index}(\mathcal{T},i)$;
@@ Appendix B Deferred Proofs
Fix a phase $i$. A new phase $i$ has just started, thus one bin of phase i1i-1 has overflowed. Since we condition on $\mathcal{O}_{i-1}$, $M_{i-1}-\frac{M_{i-1}}{\Theta(\log^{1/2}n)}$ balls have been thrown in this instance. Thus, due to the conditioning on $\mathcal{G}_{i-1}$, we also get that each bin of the instance has received at least $C_{i-1}-\frac{C_{i-1}}{\Theta(\log^{1/3}n)}$ balls. Since, the number of empty cells in each sub-subarray of $A_{i-1}$ is less than the remaining capacity of the bin we get that $\forall j,\ \texttt{empty}\left(a^{(i-1)}_{j}\right)<\frac{C}{\Theta(\log^{1/3}n)}$. Thus, we get that:
@@
after $M+\frac{M}{\Theta(\log^{1/2}n)}<M_{i}-\frac{M_{i}}{\Theta(\log^{1/2}n)}$ balls are thrown each bin is full, thus has at leat $C=C_{i}-\frac{C_{i}}{\Theta(\log^{1/3}n)}$ with probability at least $1-K_{i}/n^{2}$. Thus, for $\mathcal{E}$ it holds that:
@@
$o(v_{1},\dots,v_{d})\;=\;1+\sum_{i=1}^{d}\Big(\sigma_{i}(v_{i+1},\dots,v_{d})\cdot(v_{i}-1)+(1-\sigma_{i}(v_{i+1},\dots,v_{d}))\cdot(n_{i}-v_{i})\Big)\cdot\prod_{j<i}n_{j},$ (11)
@@ B.8 More than $\ell$ Dimensions
where $\mu_{d}:=\mathbb{E}\!\left[\|X-Y\|\right]$ for $X,Y\sim\mathcal{U}([0,1]^{d})$ independent. Anderssen et al. [ABDM76] showed that

### v3

@@ 18    $x\leftarrow\texttt{receive\_sample}(\mathcal{D})$;
19    $k_{1}\leftarrow\texttt{index}(\mathcal{T},i-1,x)$, $k_{2}\leftarrow\texttt{index}(\mathcal{T},i,x)$;
@@ Appendix B Deferred Proofs
Fix a phase $i$. A new phase $i$ has just started, thus one bin of phase i1i-1 has overflowed. Since we condition on $\mathcal{O}_{i-1}$, $M_{i-1}-\frac{M_{i-1}}{\Theta(\log^{1/2}n)}$ balls have been thrown in this instance. Thus, due to the conditioning on $\mathcal{G}_{i-1}$, we also get that each bin of the instance has received at least $C_{i-1}-\frac{C_{i-1}}{\Theta(\log^{1/3}n)}$ balls. Since, the number of empty cells in each sub-subarray of $A_{i-1}$ is less than the remaining capacity of the bin we get that $\forall j,\ \texttt{empty}\left(a^{(i-1)}_{j}\right)<\frac{C_{i-1}}{\Theta(\log^{1/3}n)}$. Thus, we get that:
@@
after $M+\frac{M}{\Theta(\log^{1/2}n)}<M_{i}-\frac{M_{i}}{\Theta(\log^{1/2}n)}$ balls are thrown each bin is full, thus has at least $C=C_{i}-\frac{C_{i}}{\Theta(\log^{1/3}n)}$ with probability at least $1-K_{i}/n^{2}$. Thus, for $\mathcal{E}$ it holds that:
@@
$o(v_{1},\dots,v_{d})\;=\;1+\sum_{i=1}^{d}\Big(\sigma_{i}(v_{i+1},\dots,v_{d})\cdot(v_{i}-1)+(1-\sigma_{i}(v_{i+1},\dots,v_{d}))\cdot(n_{i}-v_{i})\Big)\cdot\prod_{j<i}n_{j},$
@@ B.8 More than $\ell$ Dimensions
where $\mu_{d}:=\mathbb{E}\!\left[\|X-Y\|\right]$ for $X,Y\sim\mathcal{U}([0,1]^{d})$ independent. Anderssen et al. [ABD+76] showed that